Fisk Corporation is trying to improve its inventory control system and has installed an online computer at its retail stores. Fisk anticipates sales of 49,000 units per year, an ordering cost of $8 per order, and carrying costs of $1,60 per unit


 a. What is the economic ordering quantity? 

    Economic ordering quantity = _______ units 

 b. How many orders will be placed during the year? 

    orders Number of = _______ orders 

 c. What will the average inventory be? 

    Average inventory = _______ units 

 d. What is the total cost of ordering and carrying inventory? 

    Total cost = _______ 
Answer #1

EOQ=

D - annual demand, in units B ordering costs C carrying costs of the inventory (expressed as annual %) I dollar value of the inventory, per unit 2 x DxB EOQ I x C

a EOQ=Square root of ((2*49000*8)/1.60)=700 Units

b Number of Order or Annual Demand/EOQ =49000/700=70 order

total Cost of Ordering and Carrying inventory

= (order quantity/2) x holding cost per unit per year + (annual demand/order quantity) x cost per order

=700/2*1.60+49000/700*8

=$1120